# VSC-RL: Advancing Autonomous Vision-Language Agents with Variational Subgoal-Conditioned Reinforcement Learning

### v1

@@ 4 Our Approach: VSC-RL
In this section, we present our approach, Variational Subgoal-Conditioned Reinforcement Learning (VSC-RL) for enhancing vision-language agents. First, we formulate the sequential decision-making task as the variational goal-conditioned RL problem (Section 4.1). Next, we derive the new subgoal-conditioned optimization objective, SGC-ELBO consisting of (a) maximizing the subgoal-conditioned RL return (Proposition 4.1) and (b) minimizing the subgoal-conditioned difference (Proposition 4.2). We also theoretically show the equivalence between the original and derived objectives, ensuring both improved efficiency and performance guarantees. In Section 4.3, we demonstrate that VLM can effectively generate feasible subgoals form the complex goal for VSC-RL. The practical implementation is illustrated in Section 4.4. We present the overall pipeline of VSC-RL in Figure 1, and the pseudo-code of VSC-RL is summarized in Algorithm 1.
@@
     Maximize Subgoal-conditioned RL Return via Equation 5 and Equation 6
@@
  Output: $\pi$
4.1 Problem Formulation
We first formalize the real-world sequential decision-making as the variational goal-conditioned RL problem for vision-language agents. In this context, similar to Equation 1, our objective is finding goal-conditioned policy $\pi$ with the highest log evidence: $\max_{\pi}\log p_{\pi}(O|g)$ for a given goal $g$. Then, we can derive the Goal-Conditioned ELBO (GC-ELBO) of $\log p_{\pi}(O|\tau,g)$ as follows:
$\underbrace{\mathop{\mathbb{E}}_{\tau\sim p_{\pi}(\tau|g)}\left[\log p(O|\tau,g)\right]-\text{KL}(p_{\pi}(\tau|g)||p_{\pi_{\text{ref}}}(\tau|g))}_{\text{GC-ELBO}(\pi,\pi_{\text{ref}},g)},$ (2)
@@ 4.2.1 Subgoal-conditioned RL Equivalence
In the above proposition, we have successfully transformed the original goal-wise objective into the subgoal-conditioned objective which is composed of $N$ subgoals with corresponding shorter sub-trajectories. Thus, the agent can learn from the reward signals from the subgoals, thus effectively improving the learning efficiency (Jiang & Agarwal, 2018).
4.2.2 Subgoal-conditioned Difference Bound
@@
Given goal-conditioned reference policy $\pi_{\text{ref}}$ and subgoal-conditioned target policy $\pi$, the goal-conditioned KL divergence of a given goal $g$ has the upper bound of subgoal-conditioned KL divergence of corresponding subgoals $\{sg_{i}\}_{i=1}^{N}$ as followed:
@@
Therefore, from Proposition 4.2, we can directly minimize the $N$ subgoal-conditioned KL divergences which is the upper bound of the goal-conditioned KL divergence.
4.2.3 Subgoal-Conditioned ELBO
@@
$\underbrace{\mathop{\mathbb{E}}_{\tau_{i}\sim p_{\pi}(\tau_{i}|sg_{i})}\left[\log p(O|\tau_{i},sg_{i})\right]-\text{KL}(p_{\pi}(\tau_{i}|sg_{i})||p_{\pi_{\text{ref}}}(\tau_{i}|g))}_{\text{SGC-ELBO}(\pi,\pi_{\text{ref}},sg_{i},g)}.$ (3)
Equation 3 consisting of two separate learning sub-objectives: (a) maximizing the subgoal of the target policy $\pi$ and (b) minimizing the subgoal-conditioned difference with the reference policy $\pi_{\text{ref}}$. Therefore, the target agent can directly learn to resolve the subgoal $sg_{i}$ requiring a much shorter horizon requirement $g$, effectively improving the learning efficiency. Additionally, we also demonstrate the equivalence between GC-ELBO (Equation 2)and SGC-ELBO (Equation 3), thus improving learning efficiency without compromising performance guarantees.
Figure 2: Autonomous vision-language subgoal generation in AitW task. The vision-language model autonomously decomposes the goal of the complicated mobile device control task into easily achievable subgoals.
4.3 Autonomous Vision-Language Subgoal Generation
For real-world complex decision-making, it is challenging to handcraft and design the subgoals for each goal manually. VLM has exhibited a unique reasoning ability in image captioning, visual question answering, and multimodal reasoning via integrating and interpreting visual and textual information to derive meaningful insights for vision-language agents. Therefore, we use VLM as the subgoal generator which autonomously decomposes the given goal $g$ into the feasible subgoals $\{sg_{i}\}_{i=1}^{N}$. As demonstrated in the AitW task example (Figure 2), VLM can autonomously decompose the goal: ”What’s the US dollar exchange rate against the Euro?” into more specific and easily solvable subgoals including ”Open a browser”, ”Search for ”US dollar to Euro exchange rate”, and ”View the exchange rate”. In Section 5, we empirically validate that the advanced VLM can successfully decompose the goals from the subset of AitW tasks to feasible subgoals with $100\%$ success rate via human verification. Therefore, we can tell that the VLM can serve as the subgoal generator in the general case. Additionally, we also provide the example of subgoal generation in Appendix C. In the context of VLM as the autonomous vision-language subgoal generator, the optimization objective (Equation 3) can be written as
@@ 4.3 Autonomous Vision-Language Subgoal Generation
where subgoals $\{sg_{i}\}_{i=1}^{N}$ are generated by a VLM through prompting with the original goal $g$. Specifically, the pipeline of our VSC-RL is summarized in the Figure 1.
@@ 4.4 Practical Implementation
As a unified RL-based agent framework, most existing RL-based methods can easily be embedded in VSC-RL. In this paper, we mainly consider the mobile device control task to evaluate the VSC-RL, a representative real-world challenging sequential decision-making task which has dramatically drawn attention recently. Specifically, the reference agent $\pi_{\text{ref}}$ and target agent $\pi$ are both initialized as the AutoUI-Base agent (Zhang & Zhang, 2023) which is pre-trained on the Android in the Wild (AitW) datasets. To maximize the subgoal-conditioned RL objective in Equation 3, VSC-RL uses the Advantage-Weighted Regression (AWR) algorithm (Peng et al., 2019) modified by DigiRL (Bai et al., 2024) as follows:
@@
where $R_{i}$ is the binary return evaluated by the VLM (Pan et al., 2024b) and $V(s,a,sg_{i})$ is the subgoal-conditioned value function. VSC-RL minimize the subgoal-conditioned KL divergence in Equation 3 via imitation loss as follows:
@@
where $a_{\text{ref}}$ is the reference action. Similar to DigiRL (Bai et al., 2024), VSC-RL also additionally learns the instruction-level value function for filtering the sub-trajectories and accelerating the learning process.
VSC-RL adopts Gemini-1.5-Pro (Team et al., 2024) as the subgoal generator. Specifically, we in-context prompt the VLM to generate the subgoals for a given goal including human demonstration as examples. The prompt example is provided in Appendix C. Overall, the pseudo-code of VSC-RL is summarized in Algorithm 1.
@@ 5 Experiments
In this section, we first show that our VSC-RL can effectively improve learning efficiency and performance, on the toy vision-language tasks. Then, we mainly demonstrate that our VSC-RL can achieve better sample efficiency and a higher success rate than various state-of-the-art (SOTA) agents in addressing challenging and complex mobile device control tasks. We also evaluate the generalization ability of VSC-RL on unseen mobile device control tasks. We also investigate the improvement and verification of the VLM-based subgoal generator of the VSC-RL.
@@ 5.1 Experimental Settings
We first evaluate our VSC-RL on the toy vision-language tasks, MiniGrid (Chevalier-Boisvert et al., 2023). For the more complex and challenging problem, we mainly consider AitW General and Web Shopping tasks (Rawles et al., 2024), two kinds of the most challenging device control tasks for evaluation. The horizon of General and Web Shopping tasks are set to 10 and 20 steps, respectively. The success of the task is autonomously evaluated by the Gemini-1.5-Pro (Team et al., 2024) via the in-context prompting approach.
@@
For the MiniGrid, we select the PPO (Schulman et al., 2017) as the baseline, and we apply VSC-RL in the PPO for a fair comparison. For the mobile device control tasks, we compare our VSC-RL with various SOTA baselines including prompting-based agents (Set-of-Marks (Yang et al., 2023b) and AppAgent (Zhang et al., 2023)), imitation-based agents (AutoUI (Zhang & Zhang, 2023), CogAgent (Hong et al., 2024) and Filtered BC (Pan et al., 2024b)) and RL-based agents (DigiRL (Bai et al., 2024)). In MiniGrid, each method is evaluated across 5 random seeds. For AitW tasks, each method is tested on $3$ independent runs, consistent with existing works.
@@ 5.2 Experimental Results and Analysis
Table 1: The evaluated performance on the train and test datasets of the General and Web Shopping tasks. The best performance is in bold.
@@
Toy vision-language tasks on MultiRoom.
Overall, as shown in Figure 5, our VSC-RL outperforms the baseline in all tasks remarkably, especially in the difficult task with the increasing number of rooms. From the result of MultiRoom-N2-v0 shown in LABEL:fig:multiroom_2, we can tell that although PPO and VSC-RL both successfully reach $100\%$ success rate, our VSC-RL shows a better sample efficiency. For MultiRoom-N4-v0 (LABEL:fig:multiroom_4) and MultiRoom-N6-v0 (LABEL:fig:multiroom_6) where PPO is not able to learn any useful policy, while VSC-RL exhibits strong performance of $100\%$ and $80\%$ success rate respectively.
General and Web Shopping.
@@
The learning curves of AitW General and Web Shopping are summarized in Figure 7. Overall, our VSC-RL outperforms other baselines significantly in both General and Web Shopping tasks. The RL-based agents (DigiRL and VSC-RL) both show leading performance in the AitW General task. After reaching a similar performance of 0.65 success rate with DigiRL in 250 trajectories, our VSC-RL outperforms all baselines significantly, arriving at the best final performance of 0.75 success rate. Similarly, RL-based agents (DigiRL and VSC-RL) dominate all other types of agents remarkably in the Web Shopping task. Specifically, our VSC-RL can finally achieve around 0.6 success rate significantly outperforming 0.5 success rate of DigiRL.
Generalization Evaluation.
We also evaluate the generalization of our VSC-RL on train and test datasets including a range of unseen tasks, respectively. The results summarized in Section 5.2 tell us that our VSC-RL shows significant superiors in both train and test datasets. Especially, in the general tasks, VSC-RL performs approximately $13.9\%$ and $7.7\%$ better than the second best baseline on the train and test datasets, respectively. Similarly, VSC-RL also achieve the best performance on both the train and test datasets of web shopping tasks. Overall, VSC-RL can exhibit consistent performance on new tasks, showing remarkable generalization ability.
Improvement of Subgoal Generator.
We investigate the importance of the subgoal generator in our VSC-RL on the Web Shopping subsets with different horizon lengths (short, medium and long). We implement VSC-RL with the original goal instead of the subgoals generated from VLM. As shown in Figure 10, the subgoal generator can effectively improve the performance across all types of Web Shopping tasks via autonomously decomposing the original goal to the subgoals. Especially, the subgoal generator can effectively enhance the $50\%$ and $32\%$ performance in the Web Shopping medium and long tasks, respectively.
Verification of Subgoal Generator.
To investigate the quality and feasibility of the generated subgoals, we manually verify the results of subgoal generator on $135$ trajectories from the AitW human demonstration. There are $135(100\%)$ goals are decomposed into feasible subgoals successfully, and the final goal can be accomplished by reaching these subgoals sequentially. Specifically, there are $123(91.1\%)$ goals that are decomposed into subgoals completely aligning with the human demonstration. For the remaining $12(8.9\%)$ goals, the subgoal generator provides alternative subgoals different from human demonstration, but still can successfully arrive at the final goal.
@@ 5.3 Limitations and Challenges
We have empirically demonstrated that our VSC-RL can effectively address the learning efficiency issue commonly existing in complex sequential decision-making tasks. However, there still exists some limitations and challenges in VSC-RL as discussed as follows.
@@
Additionally, the VLM in VSC-RL can not only be viewed as the subgoal generator, but also as the high-level policy in the context of the hierarchical RL. It is valuable to investigate how to enhance VSC-RL with the advanced hierarchical RL approaches.
@@
In this work, we mainly consider the mobile device control task, a representative complex control problem as the evaluation benchmark. The theoretical and empirical results presented in this work imply that VSC-RL has great potential in addressing other challenging open problems such as embodied agent learning and real robotics learning.
@@ 6 Conclusion
This work investigates vision-language agents in resolving real-world complex sequential decision-making tasks. Existing promising RL-based agents often suffer from the learning efficiency issue in solving tasks with complicated goals and sparse reward signals. To address this issue, we propose VSC-RL, which can autonomously decompose the goal to subgoals and resolve them efficiently. Formulating the sequential decision-making task as a variational subgoal-conditioned RL problem with the optimization objective of SGC-ELBO. We also provide a theoretical guarantee of the performance of VSC-RL. In various benchmarks, especially in challenging mobile device control tasks, we empirically show that VSC-RL exhibits significant performance improvement and learning efficiency, remarkably outperforming existing methods.

### v2

@@ 4 Our Approach: VSC-RL
In this section, we present our approach, Variational Subgoal-Conditioned Reinforcement Learning (VSC-RL) for enhancing VLM agents in solving real-world decision-making tasks. First, we formulate the sequential decision-making task as the variational goal-conditioned RL problem (Section 4.1). Next, we derive the new subgoal-conditioned optimization objective, SGC-ELBO, consisting of (a) maximizing the subgoal-conditioned return (Proposition 4.1) and (b) minimizing the subgoal-conditioned difference (Proposition 4.2). We also theoretically show the derivation of new optimization objective, ensuring both improved learning efficiency and performance guarantees. In Section 4.3, we demonstrate that VLMs can effectively generate feasible subgoals from the complex goal for VSC-RL. The practical implementation is illustrated in Section 4.4. We present the overall pipeline of VSC-RL in Figure 1, and the pseudo-code of VSC-RL is summarised in Algorithm 1.
4.1 Problem Formulation
We first formulate the sequential decision-making as the variational goal-conditioned RL problem. In this context, similar to Equation 1, the objective is to find a goal-conditioned policy $\pi$ with the highest log evidence: $\max_{\pi}\log p_{\pi}(O|g)$ for a given goal $g$. Then, we have the Goal-Conditioned ELBO (GC-ELBO) of $\log p_{\pi}(O|\tau,g)$ as follows:
[…]
@@ 4.1 Problem Formulation
     Maximize Subgoal-conditioned Return via Equation 5 and Equation 6
@@
  Output: updated policy $\pi$
@@ 4.2 Variational Subgoal-Conditioned RL
In the above proposition, the goal-wise objective has been transformed into the subgoal-conditioned objective, which is composed of $N$ subgoals with corresponding shorter horizons. Thus, the agent can learn from these reward signals from the subgoals, thus effectively improving the learning efficiency [13].
@@
Given goal-conditioned reference policy $\pi_{\text{ref}}$ and subgoal-conditioned target policy $\pi$, the goal-conditioned KL divergence of a given goal $g$ has the upper bound of subgoal-conditioned KL divergence of corresponding subgoals $\{sg_{i}\}_{i=1}^{N}$ as follows:
@@
Therefore, from Proposition 4.2, we can directly minimize the $N$ subgoal-conditioned KL divergences, which is the upper bound of the goal-conditioned KL divergence.
Based on Proposition 4.1 and Proposition 4.2, the newly-derived optimization objective of SubGoal-Conditioned ELBO (SGC-ELBO) is as follows:
@@
Equation 3 consists of two key components: (a) maximizing the subgoal-conditioned return of the target policy $\pi$ and (b) minimizing the subgoal-conditioned difference between $\pi$ and the reference policy $\pi_{\text{ref}}$. Therefore, the agent can directly learn to resolve the subgoal $sg_{i}$ with a shorter horizon requirement, effectively improving the learning efficiency. The newly derived optimization objective of SGC-ELBO (Equation 3) can improve learning efficiency without compromising performance guarantees.
4.3 Autonomous Subgoal Generation via Vision-Language Models
Figure 2: Autonomous subgoal generation in AitW task. The VLM autonomously decomposes the goal of the complicated mobile device control task into easily achievable subgoals.
For real-world complex decision-making, it is challenging to handcraft and design the subgoals for each goal manually. VLM has exhibited a unique reasoning ability in image captioning, visual question answering, and multimodal reasoning via integrating and interpreting visual and textual information to derive meaningful insights for VLM agents. Therefore, we use VLM as the subgoal generator, which autonomously decomposes the given goal $g$ into the feasible subgoals $\{sg_{i}\}_{i=1}^{N}$. As demonstrated in the AitW task example (Figure 2), VLM can autonomously decompose the goal: "What’s the US dollar exchange rate against the Euro?" into more specific and easily solvable subgoals, including "Open a browser", "Search for "US dollar to Euro exchange rate", and "View the exchange rate". Therefore, we can tell that the VLM can serve as the subgoal generator in the general case. Additionally, we also provide the example of subgoal generation in Appendix F. In the context of VLM as the autonomous vision-language subgoal generator, the optimization objective (Equation 3) can be written as
@@ 4.3 Autonomous Subgoal Generation via Vision-Language Models
where subgoals $\{sg_{i}\}_{i=1}^{N}$ are generated by a VLM through prompting with the original goal $g$.
@@ 4.4 Practical Implementation
As a unified RL-based agent framework, most existing RL-based methods can easily be embedded in VSC-RL. In this paper, we mainly consider the mobile device and web control tasks to evaluate the VSC-RL, a representative and challenging real-world decision-making task which has drawn attention recently. Specifically, the reference agent $\pi_{\text{ref}}$ and target agent $\pi$ are both initialised as the AutoUI-Base agent [42], which is pre-trained on the Android in the Wild (AitW) datasets. To maximize the subgoal-conditioned RL objective in Equation 3, VSC-RL uses the Advantage-Weighted Regression (AWR) algorithm [25] modified by DigiRL [4] as follows:
@@
where $R_{i}$ is the binary return evaluated by the VLM [23] and $V(s,a,sg_{i})$ is the subgoal-conditioned value function. VSC-RL minimizes the subgoal-conditioned KL divergence in Equation 3 via imitation loss as follows:
@@
where $a_{\text{ref}}$ is the reference action. Similar to DigiRL [4], VSC-RL additionally learns the instruction-level value function for filtering the sub-trajectories and accelerating the learning process.
VSC-RL adopts Gemini-1.5-Pro [33] as the subgoal generator. Specifically, we in-context prompt the VLM to generate the subgoals for a given goal, including human demonstration as examples. The prompt example is provided in Appendix F. Overall, the pseudo-code of VSC-RL is summarised in Algorithm 1.
@@ 5 Experiments
In this section, we empirically demonstrate that our VSC-RL can achieve better sample efficiency and a higher success rate than various state-of-the-art (SOTA) agents in various challenging benchmarks, including AitW [27] and WebArena-Lite [17]. We also present the ablation results for evaluating the key components of VSC-RL. The implementation details and hyperparameter settings are listed in Appendix A. Additionally, we present the additional experiments on MiniGrid [9] in Appendix C. Additional experiments investigating the subgoal generator in VSC-RL are presented in Appendix D.
@@ 5.1 Experimental Settings
For the complex and challenging problem, we mainly consider AitW General and Web Shopping tasks [27], two kinds of the most challenging device control tasks for evaluation. The horizons of General and Web Shopping tasks are set to 10 and 20 steps, respectively. The success of the task is autonomously evaluated by the Gemini-1.5-Pro [33] via the in-context prompting approach. We also evaluate VSC-RL on WebArena-Lite [17], a human-verified subset of the WebArena benchmark [45] containing 165 realistic web tasks across five websites. Each task involves complex HTML-based observations with 30 steps of horizon. Following WebRL [26], we adopt the pretrained outcome-supervised reward model (ORM) to autonomously evaluate the task’s success.
@@
For the AitW tasks, we compare our VSC-RL with various SOTA baselines, including prompting-based agents (Set-of-Marks [38] and AppAgent [41]), imitation-based agents (AutoUI [42], CogAgent [12] and Filtered BC [23]) and RL-based agents (DigiRL [4]). Each method is tested on $3$ independent runs, consistent with existing works [4]. For the WebArena-Lite tasks, we compare VSC-RL with several SOTA baselines adapted to web environments, including supervised fine-tuning (SFT), Filtered BC [23], and RL-based agents (AWR [25], DigiRL [4], and WebRL [26]). Specifically, to ensure a fair comparison, we remove the curriculum learning component of WebRL so that all methods are trained solely on the same task set. Each method is tested on $1$ single run, consistent with existing works [26].
@@ 5.2 Experimental Results and Analysis
Table 1: The evaluated performance on the train and test datasets of the General and Web Shopping tasks. The best performance is highlighted.
@@
AitW.
@@
The learning curves of AitW General and Web Shopping are summarised in Figure 4. Overall, our VSC-RL outperforms other baselines significantly in both the General and Web Shopping tasks. The RL-based agents (DigiRL and VSC-RL) both show leading performance in the AitW General task. After reaching a similar performance of $65.0\%$ success rate with DigiRL in 250 trajectories, our VSC-RL outperforms all baselines significantly, arriving at the best final performance of $75\%$ success rate. Similarly, RL-based agents (DigiRL and VSC-RL) dominate all other types of agents remarkably in the Web Shopping task. Specifically, our VSC-RL can finally achieve around $60.0\%$ success rate, significantly outperforming $50.0\%$ success rate of DigiRL. We also evaluate the generalization of our VSC-RL on the test datasets, including a range of unseen tasks, respectively. The results summarised in Table 1 tell us that our VSC-RL shows significant superiority in both the train and test datasets. Especially, in the general tasks, VSC-RL performs approximately $+13.9\%$ and $+7.7\%$ better than the second-best baseline on the train and test datasets, respectively. Similarly, VSC-RL achieves the best performance on both the train and test datasets of web shopping tasks. Overall, VSC-RL can exhibit consistent performance on unseen tasks, showing remarkable generalization ability.
WebArena-Lite.
As shown in  Figure 5, we present the learn curves on the WebArena-Lite. The imitation-based agents, SFT and Filtered BC achieve relatively limited performance, with success rates stagnating around $20.6\%$ and $23.0\%$, respectively. AWR achieves approximately $28.5\%$ success rate via solely leveraging the offline RL technique, which is still limited compared to the online RL methods. DigiRL and WebRL exhibit similar performance trends, both plateauing around a $31.0\%$ success rate. Our VSC-RL consistently outperforms the other methods, achieving the highest success rate of approximately $34.5\%$. Overall, our VSC-RL can achieve superior learning efficiency and final performance on the WebArena-Lite. Specifically, the additional results shown in Appendix E demonstrate that our VSC-RL can achieve the best performance across all types of tasks on the WebArena-Lite.
Ablation Results of VSC-RL.
The ablation results on the Web Shopping task, shown in  Table 2, evaluate the impact of various key components of VSC-RL. Introducing subgoals to AutoUI can yield a marginal improvement ($+1.6\%$) on the average success rate from $16.2$% to $17.8\%$, which offers limited benefit without involving the RL process. For VSC-RL, removing subgoals entirely leads to a substantial performance drop ($-7.8\%$) compared to providing limited ($50\%$ less) subgoals ($-5.2\%$). These results imply that subgoals can efficiently improve performance by providing immediate informative reward signals in the RL process. Additionally, we investigate the different optimization components of SGC-ELBO(eq. 3) in VSC-RL. Removing the policy gradient ( Equation 5) and imitation loss ( Equation 7) result in decreased performance with $-7.8\%$ and $-10.4\%$, respectively. Overall, each optimization component of VSC-RL contributes meaningfully to its effectiveness, aligning with our main statements in Section 4.2.
Figure 5: Learning curves on WebArena-Lite.
Table 2: Ablation results of VSC-RL on the Web Shopping. The best performance is highlighted.
Method Web Shopping
Train Test Average
AutoUI 14.6% 17.7% 16.2%
   w/ subgoals 16.7% 18.8% 17.8%
VSC-RL 64.0% 59.0% 61.5%
   w/o subgoals 55.2% 52.1% 53.7%
   w/ limited ($50\%$ less) subgoals 57.3% 55.2% 56.3%
   w/o policy gradient ( Equation 5) 56.3% 51.0% 53.7%
   w/o imitation loss ( Equation 7) 55.2% 46.9% 51.1%
@@ 5.3 Limitations and Challenges
We have empirically demonstrated that our VSC-RL can effectively address the learning efficiency issue commonly existing in complex sequential decision-making tasks. However, there are still some limitations and challenges in VSC-RL, as discussed below.
@@
Additionally, the VLM in VSC-RL can not only be viewed as the subgoal generator, but also as the high-level policy in the context of hierarchical RL. It is valuable to investigate enhancing VSC-RL with the hierarchical RL approaches [43, 46].
@@
In this work, we mainly consider the mobile device and web control tasks, two representative complex control problems, as the evaluation benchmarks. The theoretical and empirical results presented in this work imply that VSC-RL has great potential in addressing other challenging open problems, such as embodied agent learning and robotics learning.
@@ 6 Conclusion
This work investigates advancing VLM agents in resolving real-world complex sequential decision-making tasks. Existing promising RL-based agents often suffer from the learning efficiency issue in solving tasks with complicated goals and sparse reward signals. To address this issue, we propose VSC-RL, which can autonomously decompose the goal to subgoals and resolve them efficiently. VSC-RL reformulates the decision-making task as a variational subgoal-conditioned RL problem with the new derived optimization objective of SGC-ELBO, thus effectively improving the learning efficiency without comprising the performance guarantee. In various benchmarks, especially in challenging mobile device and web control tasks, we empirically show that VSC-RL exhibits significant performance improvement and learning efficiency, remarkably outperforming existing methods.